The rate constant for this first-order reaction is 0.230 s 1 at 400 oC. A 3 products How long (in seconds) would it take for the concentration of A to decrease from 0.600 M to 0.230 M? Number 4.52
Answer #1

here:

[A] = 0.230 M

[A]o = 0.600 M

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln(0.230) = ln(0.600) - 0.230*t

-1.470 = -0.511 - 0.230*t

0.230*t = 0.959

t = 4.169 s

Answer: 4.169 s